OpenStudy (anonymous):

Find the equation of the parabola with focus (5, 0) and directrix x = -5.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

again we need a picture this one will open to the right, so that the \(y\) term will be squared

OpenStudy (anonymous):

ok. how do we know it will be opened to the right ?

OpenStudy (anonymous):

because \(x=-5\) is a VERTICAL line and \((5,0)\) is to the right of it

OpenStudy (anonymous):



OpenStudy (anonymous):

oh ok . whats next ?

OpenStudy (anonymous):



OpenStudy (anonymous):

we know the vertex is half way between \((5,0)\) and \(x=-5\) namely at \((0,0)\)

OpenStudy (anonymous):

and we know the distance between the vertex and the focus is \(5\) all these should be clear without using any formulas if they are not, let me know

OpenStudy (anonymous):

then since the \(y\) term is squared and the vertex is \((0,0)\) your equation is \[4p(x-h)=(y-k)^2\] or \[4\times 5(x-0)=(y-0)^2\] or just \[20x=y^2\]

OpenStudy (anonymous):

ok . thats fairly easy to understand .